# Robust fuzzy clustering with cellwise outliers

### v1

@@ 4.1 Body fat data
Table 1: List of variables of the body fat data set
ID Name Unit ID Name Unit
@@ 6 Thigh circumference (thigh) cm
Figure 9: Body fat data: analysis of the eleven variables selected
The data on body fat, available in the R package UsingR, contain physiological measurements of $250$ men – units 172 and 182 have been discarded due to erroneous values. Among the $18$ variables, we consider only those directly observed. Additionally, we exclude the variable Age, as it masks the clustering structure due to its cross effect on body phenotypes; the variables Weight and Height, as they define BMI (i.e., Body Mass Index); and the variable Fat Free Weight, since it is derived from estimated quantities not directly observed. Table 1 lists the eleven variables used for the analysis. Due to their different scales, the variables are standardized using a robust procedure that replaces the mean with the median and the standard deviation with the median absolute deviation. The presence of outlying values can be observed both by examining the boxplot (Figure LABEL:fig:_BF_box), which provides a univariate inspection, and the pair plot (Figure LABEL:fig:_BF_pair), where non-marginal outliers are visible. cellFCLUST is particularly suitable for analyzing this data set for two main reasons. First, since the variables are highly correlated, there is more information available to impute the potentially outlying values. Second, although the elongated shape of the data suggests the presence of a single cluster, this is not reasonable in this context as it would merge individuals with very different physical characteristics. This motivates considering a small value of $c$, similarly to the reasoning in [26] for the social stratification analysis. Assuming $c=2$ (see also [27]), we search for more spherical clusters that can be very close to each other up to overlapping, making the fuzzy clustering approach extremely useful.
Figure 10: Body fat data: objective function curves
@@
As a preliminary analysis, we consider the $\%$ HA and $\%$ WA obtained by varying the scale factor. Fixing $m=1.7$, which proves to be a suitable value for the degree of fuzzification in this data set (see the Supplementary Material), we choose $S=2$, as it yields $43\%$ HA and $32\%$ WA. This results to be the best choice since a small value of $S$, i.e. $S=1$, excessively fuzzifies all observations, while higher values of $S$ correspond to an overly high percentage of hard assignments given the data configuration – for $S=3$ and $S=4$, the percentages of HA are $89\%$ and $98\%$, while those of WA are $4\%$ and $1\%$, respectively. After selecting the fuzzifier parameter, the scale factor, and the constant for the eigenvalue ratio, we determine the values of $K$ and $\alpha$ based on the objective function curves reported in Figure 10. Specifically, we vary $K\in\{1,\ldots,8\}$ and $\alpha$ from the set that increases by $2.5\%$ increments from $2.5\%$ to $20\%$, including also $0\%$ and $1\%$. Looking at Figure 10, we can choose $K=4$, since it is the value from which the increase in the objective function does not justify an increase in the number of clusters. However, there is no clear indication of the optimal contamination level, which can be further investigated via the plots of the knee points. Following the same reasoning illustrated for the artificial data, we select $\alpha=0.05$, as this is the level at which the median difference between the knee points of the $\Delta$ values and the corresponding contamination level is close to zero, and the variability is minimal (Figure 11). This choice can be corroborated via the inspection of the $\Delta$ plots provided in the Supplementary Material.
@@
Figure 11: Body fat data: difference between the knee point of $\Delta_{ij}$ and the $\alpha$ value. The curve shows the median differences across variables as $\alpha$ varies, with $K=2$. The shaded gray area represents the variability
@@
Figure 12: Body fat data: clustering results (Cluster 1: black, Cluster 2: blue, Cluster 3: red, Cluster 4: green). (a) Pair plot of the data; crosses indicate outlying values in at least one of the two variables. (b) Tetrahedron plot showing cluster assignments; color intensity and point position reflect the degree of fuzziness (i.e., hard and soft memberships)
@@
The clustering results from cellFCLUST, using the tuning parameters set as previously detailed, are shown in Figure 12. It includes a tetrahedral representation (Figure LABEL:fig:_BF_Fuzzy), where points closer to the vertices have higher membership to the corresponding cluster, while those on the edges, or generally among vertices, are more fuzzified. The four clusters can be interpreted based on their configuration across variables (Figure LABEL:fig:_BF_PairLab). Cluster 1 (black) consists of individuals with the lowest measurements for all variables, including BMI, whose values, on the original scale, fall within the normal weight range, i.e. BMI <25<25. Cluster 2 (blue) includes men with generally higher measurements than Cluster 1 across the eleven variables, with BMI values still mostly within the normal range. For Cluster 3 (red), some overlap with Cluster 2 is observed on specific variables (e.g., wrist), while BMI predominantly falls within the overweight range, i.e. BMI $\in[25,30)$. Finally, Cluster 4 (green) groups individuals in the severe overweight to obese range – the latter corresponds to BMI $\geq 30$ – with the highest values for all variables. As expected, Cluster 1 shows the highest percentage of hard assignments ($60\%$ of the units in the cluster), and only $4$ observations with weak assignments. The fuzziness of these observations is directed toward Cluster 2, with membership degrees to the latter ranging from $0.19$ to $0.37$. Cluster 2, which has $52\%$ HA, contains $31$ observations with maximum membership below $0.9$; among them, $24$ have their second-highest membership in Cluster 3, and $7$ in Cluster 1. Cluster 3 includes $23$ weak assignments, with $13$ represented by men whose second-highest membership lies in Cluster 2. The remaining $10$ would be secondarily assigned to Cluster 4, which, in turn, contains $17$ weakly assigned units whose second-highest membership is in Cluster 3. These individuals may represent a subgroup of men for whom further analyses could be conducted – for instance, to assess whether they should follow specific dietary plans designed for individuals with obesity.
@@
Table 2: Proportion of outlying values detected by cellFCLUST in the body fat data set per variable and cluster. Total per row results in a fixed contamination level, which is $0.048$ in this case since nh=n-h= $250-\left\lceil 0.95\times 250\right\rceil=250-238=12$ observations
@@ 4.2 Well-being: OECD regional data
Table 3: Number of regions by OECD country
@@
Figure 13: OECD data: proportion of hard assignments depending on $K$ and $\alpha$ for four different levels of $S$
Unlike the previous application, we choose a higher value of $c$, allowing for elongated clusters. This choice is motivated by the reasonable assumption of correlation between variables (e.g., Education and Income) within the same cluster, as well as different degrees of variability between clusters. Specifically, we impose $c=50$, which corresponds to the order of magnitude of the ratio between the eigenvalues of the covariance matrix computed on the entire population. Since fuzzy models are scale-dependent and a constraint on the eigenvalues is imposed, careful consideration should be given to the preprocessing of the variables. Given the nature of the indicators, we consider a selected number of scaling factors $S=\{1,2,5,10\}$, which allow to adjust the percentage of hard assignments while maintaining high comprehensibility of the scaled scores. We select $S=5$ as the percentage of HA approaches $74\%$ – it ranges from $14\%$ to $91\%$ depending on the choices of $K$ and $\alpha$, as shown in Figure LABEL:fig:_OECD_S5. The fuzzifier parameter $m$ is set to $1.8$ as it provides desirable levels of WA, with $29$ weakly assigned regions out of $447$ ($18$ for $m=1.6$ and $55$ for $m=2$, as shown in the Supplementary Material). Once defined $c$, $S$, and $m$, it is possible to compute the objective function curves (Figure 14). Looking at their behavior as $K$ and $\alpha$ vary, a sensible choice for $K$ results to be $5$. Indeed, the objective function value shows substantial jumps for $K=1,\ldots,5$ across the $\alpha$ levels. After $K=5$, the values become closer to each other when $\alpha$ is sufficiently high, that is, from $7.5\%$ onward, providing evidence of this level of contamination. To support this choice, we provide also the plot of the knee points in Figure 15.
Figure 14: OECD data: objective function curves
Figure 15: OECD data: difference between the knee point of $\Delta_{ij}$ and the $\alpha$ value. The curve shows the median differences across variables as $\alpha$ varies, with $K=5$. The shaded gray area represents the variability
@@
Figure 17: OECD data: outliers of selected regions (yellow: reliable cells; blue: contaminated cells imputed with a higher value than the original one; red: contaminated cells imputed with a lower value than the original one; white: missing values)
@@ 5 Conclusions
We have introduced a new methodology, called cellFCLUST, for fuzzy clustering and cellwise outlier detection. Following the recent paradigm of cellwise contamination, which considers single cells of a data matrix to be potentially contaminated, we have developed an EM algorithm in which these cells are flagged, treated as missing values, and imputed rather than discarded, before estimating the model parameters. If not properly identified, outlying values and units with a doubtful assignment can interplay in ways that change the clustering structure and bias the parameter estimates.
Through two real-world applications, we have shown the effectiveness and usefulness of the proposed model. The first one focuses on analyzing risk levels among individuals in the overweight to obese range, leveraging fuzzification to identify observations whose measurements warrant closer inspection. The second application concerns indicators used to study well-being across regions of the OECD countries. In this framework, cellFCLUST enables the identification of common patterns among countries in terms of well-being, serving two main purposes: on one hand, to reveal different behaviors among regions within the same country – for example, California is grouped with Northern European countries based on its measurement; and on the other hand, to highlight indicators on which some countries (and their regions) display anomalous values relative to the cluster they belong to. This is the case, for instance, of Switzerland, whose income is higher even than that of Northern European countries.
A crucial role in the implementation of cellFCLUST is played by its tuning parameters, as the clustering and outlier detection results are sensitive to their choice. We have illustrated their effects on artificial data, providing guidance to users for their selection. When prior information on the real data under study is available, we suggest using it to tune the corresponding parameters. It is worth noting that all these parameters – namely, the scale factor $S$, the fuzzifier parameter $m$, the constant $c$ for the eigenvalue-ratio constraint, the number of clusters $K$, and the contamination level $\alpha$ – are interconnected and require preliminary analyses for appropriate setting. The dependence of cellFCLUST on these parameters offers users a flexible methodology that accommodates on the purpose of the analysis and prevents issues in the parameter estimation. However, selecting suitable values may require in-depth examination of the data before applying the steps described at the end of Section 3. This remains an open issue, leaving room for improvement and future developments aimed at identifying a procedure for the simultaneous selection of all tuning parameters in cellFCLUST.
@@
Supplementary Material to “Robust fuzzy clustering with cellwise outliers”
[…]
6 Additional results on the effects of the tuning parameters
In Figure 18, we display the $\Delta_{ij}$ values with $\alpha=0.05$ for the last three variables of the artificial data set, whose generation is detailed in the main article (first example). As also for the first two variables, the choice of $\alpha=0.05$, which corresponds to the true level of contamination in the data, is confirmed by the $\Delta$ plots.
Figure 18: Artificial data: $\Delta$ plot for the last three variables of the data set, where units are sorted according to their $\Delta$ values. Vertical, red line corresponds to $\alpha=0.05$ when $K=2$
The discussion on the effect of the constant $c$ for the eigenvalue-ratio constraint, reported in the main article, can be expanded to gather the potential misclassifications induced by a small value of $c$, especially when there is no justification for such a strong restriction – as, for instance, in Section 4.1 of the main article. We generate here two clusters similarly to the second example in Section 3 of the main article, but assume a spherical shape for the second cluster (green). When $c=50$, the clustering structure is well recovered, except for two points misclassified into the first (red) cluster (Figure LABEL:fig:_examplesup_c50). Contrarily, when $c=1$, the clusters are constrained to be spherical and of similar size, which causes a high misclassification rate only partially mitigated by a higher fuzzification (Figure LABEL:fig:_examplesup_c1). It is worth noting that in both cases $m=1.9$.
Figure 19: Artificial data: additional example on the effect of the constant $c$ for the eigenvalue-ratio constraint. Stars denote observations that have outlying values in the first variable
7 Additional results for the real data analyses
7.1 Body fat data set
Figure 20: Body fat data: tetrahedron plot showing cluster assignments (Cluster 1: black, Cluster 2: blue, Cluster 3: red, Cluster 4: green) when $m$ varies. Color intensity and point position reflect the degree of fuzziness (i.e., hard and soft memberships)
The preliminary analysis described in Section 4.1 of the main article on the body fat data set allows us to choose the fuzzifier parameter $m$. Specifically, we select $m$ by examining the fuzzification obtained by cellFCLUST. Recalling that we select $c=2$ to avoid obtaining only one elongated cluster, we expect a sensible proportion of weakly assigned units (i.e., $u_{ik}<0.9$, when observation $i$ belongs to cluster $k$) due to the closeness of the clusters. Figure 20 shows the hard and soft assignments to the clusters as $m$ varies. When $m=1.3$, very few observations are weakly assigned; the same holds for $m=1.5$. More sensible values for the fuzzifier parameter appear to be $m=1.7$ and $m=1.9$. Between these two, we choose the former since the difference lies mostly in the fuzzification between Cluster 2 (blue) and Cluster 3 (red), while the main interest is likely between the third and fourth clusters, where the latter mainly contains obese men (green cluster). It is worth highlighting that the choices of the tuning parameters are strongly related to each other. Therefore, in-depth analyses of their settings are necessary for the user when no prior information is available about the data – such as that used for setting $c$.
We also provide here the plot of the $\Delta$ values per variable (Figure 21) to corroborate the choice of $\alpha=0.05$, as discussed in the main article. As we can see in the figure, the chosen contamination level turns out to be suitable.
Figure 21: Body fat data: $\Delta$ plot for each variable, where units are sorted according to their $\Delta$ values. Vertical, red line corresponds to $\alpha=0.05$ when $K=4$
7.2 OECD regional data set on well-being
The choice of the parameter $m$ used in Section 4.2 is based on the proportions of weak assignments, which are reported in Figure 22. We consider here a reasonable subset for $G$ and $\alpha$, i.e., $\{3,\ldots,6\}$ and $\{0.05,0.075,0.1,0.125\}$, respectively, with different candidate levels for $m$. The selected value, which corresponds to $1.8$ (Figure LABEL:fig:_OECD_m18), provides a range between $5\%$ and $12\%$ of weak assignments, which translates to $22$ and $59$ regions, depending on the other parameters. We can also see that the proportion of hard assignments is relatively stable across different values of $m$. In Table 4, we report the complete list of $29$ weakly assigned regions with the respective membership across the five clusters.
In Figure 23, it is possible to inspect the $\Delta$ plot for each variable. The level of $\alpha$ proves to be sufficient in every variable supporting the chosen contamination level, which is $0.075$. Furthermore, Figure 24 shows the unreliable cells for every region grouped by cluster, where the red and blue cells represent lower and higher imputed values than the actual data, respectively, while yellow represents reliable cells and white are missing data.
Figure 22: OECD data: proportion of weak assignments depending on a subset of $K$ and $\alpha$, for five different levels of $m$.
Table 4: Weak assignments for the OECD data
Country Region 1 2 3 4 5
Cluster 1 Canada Newfoundland and Labrador 0.86 0.01 0.03 0.01 0.09
United States California 0.84 0.01 0.07 0.02 0.05
Cluster 2 Colombia San Andrés 0.20 0.43 0.08 0.20 0.10
Colombia Amazonas 0.02 0.62 0.01 0.31 0.04
Colombia Guainía 0.02 0.47 0.01 0.45 0.06
Colombia Guaviare 0.03 0.47 0.01 0.40 0.09
France Mayotte 0.05 0.75 0.04 0.14 0.03
Mexico Sonora 0.03 0.65 0.01 0.29 0.03
Cluster 4 Canada Nunavut 0.02 0.02 0.01 0.57 0.38
Colombia Cauca 0.02 0.03 0.00 0.90 0.05
Costa Rica Central 0.02 0.01 0.01 0.89 0.08
Costa Rica Central Pacific 0.01 0.02 0.00 0.54 0.43
Costa Rica Brunca 0.02 0.10 0.01 0.56 0.31
Costa Rica Huetar Caribbean 0.01 0.03 0.00 0.81 0.15
France Guadeloupe 0.14 0.26 0.04 0.33 0.24
Greece North Aegean 0.03 0.03 0.01 0.78 0.15
Greece Peloponnese 0.04 0.02 0.01 0.70 0.23
Cluster 5 Belgium Brussels Capital Region 0.09 0.01 0.02 0.03 0.86
Chile Magallanes and Chilean Antarctica 0.14 0.02 0.10 0.06 0.69
Chile Arica y Parinacota 0.09 0.02 0.02 0.09 0.79
France Corsica 0.06 0.01 0.01 0.03 0.90
France Martinique 0.32 0.05 0.04 0.21 0.38
Greece South Aegean 0.03 0.02 0.00 0.15 0.81
Greece Western Macedonia 0.03 0.03 0.00 0.24 0.70
Greece Epirus 0.01 0.01 0.00 0.39 0.58
Greece Ionian Islands 0.04 0.01 0.00 0.20 0.74
Japan Shikoku 0.31 0.16 0.01 0.02 0.50
Portugal Azores (autonomous region) 0.03 0.19 0.03 0.04 0.70
United States Hawaii 0.04 0.02 0.03 0.01 0.90
Figure 23: OECD data: $\Delta$ plot for each variable, where units are sorted according to their $\Delta$ values. Vertical, red line corresponds to $\alpha=0.075$ when $K=5$.
Figure 24: OECD data: outlying cells per cluster

### v2

@@ 5.1 Body fat data
The data on body fat, available in the R package UsingR, contain physiological measurements of $250$ men – observations 172 and 182 have been discarded due to erroneous values. Among the $18$ variables, we consider only those directly observed. Additionally, we exclude the variable Age, as it masks the clustering structure due to its cross effect on body phenotypes; the variables Weight and Height, as they define BMI (i.e., Body Mass Index); and the variable Fat Free Weight, since it is derived from estimated quantities not directly observed. Table 2 lists the eleven variables used for the analysis. Due to their different scales, the variables are standardized using a robust procedure that replaces the mean with the median and the standard deviation with the median absolute deviation. The presence of outlying values can be observed both by examining the boxplot (Figure LABEL:fig:_BF_box), which provides a univariate inspection, and the pair plot (Figure LABEL:fig:_BF_pair), where non-marginal outliers are visible. cellFCLUST is particularly suitable for analyzing this data set for two main reasons. First, since the variables are highly correlated, there is more information available to impute the potentially outlying values. Second, although the elongated shape of the data suggests the presence of a single cluster, this is not reasonable in this context as it would merge individuals with very different physical characteristics. This motivates considering a small value of $c$, similarly to the reasoning in [35] for the social stratification analysis. Assuming $c=2$, we search for more spherical clusters that can sometimes be very close to each other, up to overlapping, making the fuzzy clustering approach extremely useful.
Table 2: List of variables of the body fat data set.
ID Name Measur. unit ID Name Measur. unit
@@ 6 Thigh circumference (thigh) cm
Figure 7: Body fat data: analysis of the eleven variables selected.
Figure 8: Body fat data: objective function curves.
@@
As a preliminary analysis, we consider the proportions of HA and WA obtained by varying the scale factor. Fixing $m=1.7$, which proves to be a suitable value for the degree of fuzzification in this data set (see the Supplementary Material), we choose $S=2$, as it yields a proportion of HA and WA of 0.43 and 0.32, respectively. This results to be the best choice since a small value of $S$, i.e. $S=1$, excessively fuzzifies all observations, while higher values of $S$ correspond to an overly high proportion of hard assignments given the data configuration – for $S=3$ and $S=4$, the proportions of HA are 0.89 and 0.98, while those of WA are 0.04 and 0.01, respectively. After selecting the fuzzifier parameter, the scale factor, and the constant for the eigenvalue ratio, we determine the values of $K$ and $\alpha$ based on the objective function curves reported in Figure 8. Specifically, we vary $K\in\{1,\ldots,8\}$ and $\alpha$ from the set that increases by $0.025$ increments from $0.025$ to $0.20$, including also 0 and $0.01$. Looking at Figure 8, we can choose $K=4$, since it is the value from which the increase in the objective function does not justify an increase in the number of clusters. However, there is no clear indication of the optimal contamination level, which can be further investigated via the plots of the knee points. Following the same reasoning illustrated for the artificial data, we select $\alpha=0.05$, as this is the level at which the median difference between the knee points of the $\Delta$ values and the corresponding contamination level is close to zero, and the variability is minimal (Figure 9). This choice can be corroborated via the inspection of the $\Delta$ plots provided in the Supplementary Material.
@@
Figure 9: Body fat data: difference between the knee point of $\Delta_{ij}$ and the $\alpha$ value. The plot shows the median differences across variables as $\alpha$ varies, with $K=4$. The shaded gray area represents the variability.
@@
Figure 10: Body fat data: clustering results (Cluster 1: black, Cluster 2: blue, Cluster 3: red, Cluster 4: green). (a) Pair plot of the data; crosses indicate outlying values in at least one of the two variables. (b) Tetrahedron plot showing cluster assignments; color intensity and point position reflect the degree of fuzziness (i.e., hard and soft memberships).
@@
The clustering results from cellFCLUST, using the tuning parameters set as previously detailed, are shown in Figure 10. It includes a tetrahedral representation (Figure LABEL:fig:_BF_Fuzzy), where points closer to the vertices have higher membership to the corresponding cluster, while those on the edges, or generally among vertices, are more fuzzified. The four clusters can be interpreted based on their configuration across variables (Figure LABEL:fig:_BF_PairLab). Cluster 1 (black) consists of individuals with the lowest measurements for all variables, including BMI, whose values, on the original scale, fall within the normal weight range, i.e. BMI <25<25. Cluster 2 (blue) includes men with generally higher measurements than Cluster 1 across the eleven variables, with BMI values still mostly within the normal range. For Cluster 3 (red), some overlap with Cluster 2 is observed on specific variables (e.g., wrist), while BMI predominantly falls within the overweight range, i.e. BMI $\in[25,30)$. Finally, Cluster 4 (green) groups individuals in the severe overweight to obese range – the latter corresponds to BMI $\geq 30$ – with the highest values for all variables. As expected, Cluster 1 shows the highest proportion of HA (0.60 of the observations in the cluster), and only $4$ observations with WA. The fuzziness of these observations is directed toward Cluster 2, with membership degrees to the latter ranging from $0.19$ to $0.37$. Cluster 2, which has a proportion of HA of 0.52, contains $31$ observations with maximum membership below $0.9$; among them, $24$ have their second-highest membership in Cluster 3, and $7$ in Cluster 1. Cluster 3 includes $23$ weak assignments, with $13$ represented by men whose second-highest membership lies in Cluster 2. The remaining $10$ would be secondarily assigned to Cluster 4, which, in turn, contains $17$ weakly assigned observations whose second-highest membership is in Cluster 3. These individuals may represent a subgroup of men for whom further analyses could be conducted – for instance, to assess whether they should follow specific dietary plans designed for individuals with obesity.
@@
Table 3: Proportion of outlying values detected by cellFCLUST in the body fat data set per variable and cluster. Total per row results in a fixed contamination level, which is $0.048$ in this case since nh=n-h= $250-\left\lceil 0.95\times 250\right\rceil=250-238=12$ observations.
@@ 5.2 Well-being: OECD regional data
Table 4: Number of regions by OECD country.
@@
Figure 11: OECD data: proportion of hard assignments depending on $K$ and $\alpha$ for four different levels of $S$.
Unlike the previous application, we choose a higher value of $c$, allowing for elongated clusters. This choice is motivated by the reasonable assumption of correlation between variables (e.g., Education and Income) within the same cluster, as well as different degrees of variability between clusters. Specifically, we impose $c=50$, which corresponds to the order of magnitude of the ratio between the eigenvalues of the covariance matrix computed on the entire data set. Since fuzzy models are scale-dependent and a constraint on the eigenvalues is imposed, careful consideration should be given to the preprocessing of the variables. Given the nature of the indicators, we consider a selected number of scaling factors $S=\{1,2,5,10\}$, which allow to adjust the proportion of hard assignments while maintaining high comprehensibility of the scaled scores. We select $S=5$ as the proportion of HA approaches $0.74$ – it ranges from $0.14$ to $0.91$ depending on the choices of $K$ and $\alpha$, as shown in Figure LABEL:fig:_OECD_S5. The fuzzifier parameter $m$ is set to $1.8$ as it provides desirable levels of WA, with $29$ weakly assigned regions out of $447$ ($18$ for $m=1.6$ and $55$ for $m=2$, as shown in the Supplementary Material). Once defined $c$, $S$, and $m$, it is possible to compute the objective function curves (Figure 12). Looking at their behavior as $K$ and $\alpha$ vary, a sensible choice for $K$ results to be $5$. Indeed, the objective function value shows substantial jumps for $K=1,\ldots,5$ across the $\alpha$ levels. After $K=5$, the values become closer to each other when $\alpha$ is sufficiently high, that is, from $0.075$ onward, providing evidence of this level of contamination. To support this choice, we provide also the plot of the knee points in Figure 13.
Figure 12: OECD data: objective function curves.
Figure 13: OECD data: difference between the knee point of $\Delta_{ij}$ and the $\alpha$ value. The plot shows the median differences across variables as $\alpha$ varies, with $K=5$. The shaded gray area represents the variability.
@@
Figure 15: OECD data: outlying cells of selected regions (yellow: reliable cells; blue: detected contaminated cells imputed with a higher value than the original one; red: detected contaminated cells imputed with a lower value than the original one; white: missing values).
@@ 6 Conclusions
A new methodology, called cellFCLUST, has been introduced for fuzzy clustering and cellwise outlier detection. Following the recent paradigm of cellwise contamination, which considers single cells of a data matrix to be potentially contaminated, we have developed an algorithm intended for flagging these cellwise outliers, which are then treated as missing values, and imputed rather than discarded, before estimating the model parameters. If not properly identified, outlying values and fuzzy assignments can interplay in ways that change the clustering structure and bias the parameter estimates.
Through a simulation study and two real-world applications, we have shown the effectiveness and usefulness of the proposed model. The former illustrate the performance of cellFCLUST in cluster recovery, parameter estimation and outlier detection compared to other fuzzy clustering methodologies. The results demonstrate the advantages of the proposal when cellwise contamination occurs. Regarding the real data analyses, the first one focuses on analyzing risk levels among individuals in the overweight to obese range, leveraging fuzzification to identify observations whose measurements warrant closer inspection; the second application concerns indicators used to study well-being across regions of the OECD countries. In this framework, cellFCLUST enables the identification of common patterns among countries in terms of well-being, serving two main purposes: on one hand, to reveal different behaviors among regions within the same country – for example, California is grouped with Northern European countries based on its measurement; and on the other hand, to highlight indicators on which some countries (and their regions) display anomalous values relative to the cluster they belong to. This is the case, for instance, of Switzerland, whose income is higher even than that of Northern European countries.
A crucial role in the implementation of cellFCLUST is played by its tuning parameters, as the clustering and outlier detection results can be sensitive to their choice. We have illustrated their effects on artificial data, providing guidance to users for their selection. It is worth noting that all these parameters – namely, the number of clusters $K$, the contamination level $\alpha$, the constant $c$ for the eigenvalue-ratio constraint, the fuzzifier parameter $m$, and the scale factor $S$ – are interconnected and require interrelated analyses for appropriate setting. The dependence of cellFCLUST on these parameters offers users a flexible methodology that accommodates on the purpose of the analysis and prevents issues in the parameter estimation. Simplifying these connected tasks remains an open issue, leaving room for improvement and future developments aimed at identifying a procedure for the simultaneous selection of all tuning parameters in cellFCLUST. Future work could also explore extending the robust fuzzy clustering approach with a factorial structure for the cluster covariance matrices [36] to the the cellwise contamination framework.
@@
Code availability
The R code for the implementation of cellFCLUST and the simulation study in Section 3 is available at https://github.com/giorgiazaccaria/cellFCLUST.
@@
Funding
The research of Giorgia Zaccaria and Francesca Greselin was supported by Milano-Bicocca University Fund for Scientific Research, 2023-ATE-0448. Francesca Greselin’s research was also supported by PRIN2022 - 2022LANNK
C. The research of Luis A. García-Escudero and Agustín Mayo-Íscar was partially supported by grant PID2021-128314NB-I00 funded by MCIN/AEI/10.
13039/501100011033/FEDER, UE and Junta Castilla y León grant VA064G2
4.